OpenStudy (anonymous):

i'm in math 2 and im not going to be in school tomorrow and she gave me some notes and i dont know how to do the stuff she gave me. its about circles and i dont know how to do it.

OpenStudy (amistre64):

What about circles? C=2pi*r, A=pi*r^2, angles?, equations for circles such as: (x+h)^2 + (y+k)^2 = r^2?

OpenStudy (anonymous):

like where chord radius diameter tangent arc magor arc minor arc semicircle and secant is

OpenStudy (amistre64):

Chord means: a line that cross a circle an touches at two points. Diameter means: how wide the circle is across. Radius means: half of its diameter. A tangent is: a line that touches the cirlce at only one point. Arcs are just parts of the outside of a circle. A semicircle tends to be just half of the circle. Major and minor and secant I am not familiar with.

OpenStudy (anonymous):

ok thx